# Halfway to the Peak: The JWST MIRI 5.6 micron number counts and source population

### v1

@@ Abstract
We present an analysis of 8 JWST Mid-Infrared Instrument 5.6 micron images with depths of 0.1Jy. We detect 2854 sources within our combined area of 18.4 sq.arcmin – a increase in source density over earlier IRAC channel 3 data. We compute the MIRI 5.6 m number counts including an analysis of the field-to-field variation. Relative to earlier published MIRI 5.6 m counts, our counts have a more pronounced knee, at roughly 2 Jy. The location and amplitude of the counts at the knee are consistent with the Cowley et al. (2018) model predictions, although these models tend to overpredict the counts below the knee. In areas of overlap, 84 % of the MIRI sources have a counterpart in the COSMOS2020 catalog. These MIRI sources have redshifts that are mostly in the , with a tail out to . They are predominantly moderate to low stellar masses (M⊙) main sequence star-forming galaxies suggesting that with 2 hr exposures, MIRI can reach well below at cosmic noon and reach higher mass systems out to . Nearly 70% of the COSMOS2020 sources in areas of overlap now have a data point at 5.6 m (rest-frame near-IR at cosmic noon) which allows for more accurate stellar population parameter estimates. Finally, we discover 31 MIRI-bright sources not in COSMOS2020. A cross-match with IRAC channel 1 suggests that 10-20 % of these are likely lower mass (MM⊙), dusty galaxies. The rest (80–90%) are consistent with more massive, but still very dusty galaxies at .
revtex4-1.clsrevtex4-2.cls
@@ 1 Introduction
Extragalactic astronomy aims to study how galaxies form and evolve across time. To do so, we build multiwavelength galaxy surveys which allow us to compare the data with spectral energy distribution (SED) models to infer redshifts and stellar population parameters (e.g. Brammer et al., 2008; Weaver et al., 2022a; Wang et al., 2024). Mid-IR data are particularly critical to this effort since they probe the rest-frame near-IR in the cosmic noon epoch and earlier which allows for much more accurate stellar masses and star formation rates (SFRs) (e.g. Elsner et al., 2008; Stefanon et al., 2015; Muzzin2009l; Martis et al., 2023; La Torre et al., 2024). Indeed, over the past two decades Spitzer IRAC data – especially its first two channels, available even in the warm mission (Lacy et al., 2021; Annunziatella et al., 2023) have been a critical component of these multiwavelength surveys. With IRAC channel 1 and 2, we reached stellar masses of M at cosmic noon, i.e. well below the knee of the SMF (Elsner et al., 2008; Madau & Dickinson, 2014), but only detected the most massive galaxies () at (e.g. Stefanon et al., 2015). The IRAC channel 3, at 5.8 m, which critically covers the rest-frame of the stellar 1.6 m bump at was relatively underutilized because it was not available during the extended warm mission and was the least sensitive IRAC channel even during the cold mission.
@@
The JWST (Gardner et al., 2006), which launched on December 25, 2021, is already revolutionizing infrared astronomy with its unprecedented performance (Rigby et al., 2023). In particular, the Mid-InfraRed Instrument (MIRI; Rieke et al., 2015) allows for significantly greater sensitivity and angular resolution relative to the Spitzer Space Telescope. Recently published JWST/MIRI number counts at 7.7 m, 10 m and 15 m bands (Ling et al., 2022; Wu et al., 2023; Kirkpatrick et al., 2023; Stone et al., 2024) show dramatic improvement in the depth reached, even with moderate exposure times, relative to prior measurements from Spitzer and ISO. These counts have already been used in constraining galaxy and black hole evolution models (Kim et al., 2024). Number counts at the shortest MIRI wavelength (5.6 m) are much more scarse (Yang et al., 2023; Stone et al., 2024), but they are critical as this band samples the stellar 1.6 m bump at and thus is critical in testing our models of the galaxy population at cosmic noon.
@@
In addition, sampling the rest-frame near-IR at cosmic noon, the 5.6 m band is much less sensitive to the effects of dust obscuration even than traditionally ‘dust clean’ bands such as the band, which is rest-frame -band at the same redshifts. This insensitivity to dust obcuration is important as in the more than two decades since the discovery of the Cosmic Infrared Background (e.g. Puget et al., 1996) it has become abundantly clear that the bulk of star-formation activity at cosmic noon and beyond takes place in dust obscured galaxies (see Casey et al. 2014 for a review) which in their extreme are missed in UV/optical surveys (see e.g. Hughes et al., 1998; Sajina et al., 2006; Pope et al., 2008; Zavala et al., 2021, among many others). Recently, JWST data have further highlighted this by finding that even deep HST data miss the reddest/most dust obscured parts of the galaxy population (e.g. Labbe et al., 2023; Barrufet et al., 2023; Williams et al., 2023).
@@
In this paper, we use deep MIRI 5.6 m images to provide a first look at the properties of the 5.6 m number counts and source population based on images obtained with nearly 2 hr exposure times. This wavelength corresponds roughly to the rest-frame -band at cosmic noon (), thus probing primarily stellar mass at these critical redshifts. This paper is structured as follows. In Section 2 we present the MIRI 5.6 m imaging data and data reduction. In Section 3 we present the source detection and photometry and the verification of the latter. In Section 4, we show the key results of our work. These include the 5.6 m number counts (Section 4.2); the redshift distribution and other properties of the MIRI 5.6 m sources with counterparts in COSMOS2020 (Section 4.3); and lastly MIRI 5.6 m sources without counterparts in COSMOS2020 suggesting very red colors (Section 4.4). Throughout we adopt the AB magnitude system (Oke & Gunn, 1974). We adopt the cosmology model from the COSMOS2020 catalog (Weaver et al., 2022a), which is a standard CDM cosmology with H0 = 70 km/s/Mpc, = 0.3, and = 0.7.
@@ 2.1 MIRI imaging data
Our data come from JWST/MIRI imaging with the 5.6 m filter obtained as part of the GO1 program “Halfway to the Peak: A Bridge Program To Map Coeval Star Formation and Supermassive Black Hole Growth” (PIs Pope, Sajina, Yan; PID 1762; Pope et al., 2021). This program observed 8 targets with the MIRI/MRS instrument: details on the targets and MRS spectra analysis are provided in Young et al. (2023). Simultaneous with the MIRI/MRS observations, we obtained MIRI 5.6 m imaging in fields adjacent to the MRS targets. This parallel observing means our imaging fields are effectively blank fields and therefore ideal for statistical studies such as number counts. In total, we have six fields in the First Look Survey (FLS) field and two in the COSMOS field. The FLS fields were observed in July/August 2022 and the COSMOS fields were observed in December 2022.
@@
We obtain data over the MIRI imager with an FOV of 113′′ as well as the smaller MIRI coronograph with a FOV of 24′′ (as seen in Figure 1). For our eight pointings, this adds up to a total area of just under 19.4 sq.arcmin. The effective area we use is a bit more restricted after masking out noisy edges and bright stars. This is discussed in Section 3.2.
@@ 2.2 Data reduction
We first obtained our MIRI images from the archive already reduced using the JWST operations pipeline build 10.1 (science calibration pipeline version 1.12.5; CRDS version 11.17.6, context jwst_1193.pmap)111As of the time of writing, this was the latest pipeline build that affected MIRI imaging. See https://jwst-docs.stsci.edu/jwst-science-calibration-pipeline-overview/jwst-operations-pipeline-build-information. Figure 2 left shows an example of one of our images. Within these pipeline-reduced images we noted both the presence of an overall background level (the median pixel value for the FLS images was 1.23, for the COSMOS images it was 1.52). This difference in overall background level is unsurprising given that FLS is a northern field, whereas COSMOS is equatorial and as such sees a higher level of Zodiacal light, see example discussion in Sanders et al. (2007). We also noted some clear striping especially toward the bottom quarter of each image.
@@
We first attempted a simple background subtraction motivated by the fact that the stripping patterns were fairly consistent among the images. We thus constructed a median background image where we used the segmentation map provided with pipeline-processed images to mask out bright sources in the fields. To avoid any remaining source signal, we added a 5 pixel “buffer” around each source within the segmentation maps. Due to the difference in overall pixel level between the images, we rescaled each source-subtracted image to a median of one. Figure 2 middle shows an example of our image after this basic median background subtraction when we only use the FLS images for constructing the background. It is clear that this method works well in removing the vast majority of the striping though some residual structure remains due to second order differences between images. In addition, since we only have two COSMOS images we could not construct a COSMOS only median using this method. We attempted a combined median background in FLS and COSMOS; however, this led to the COSMOS fields (which have less weight in this median with 2/8 images only) to have higher noise levels. Specifically, using a single median we found 5 levels of 0.095 Jy in FLS and 0.135 Jy in COSMOS. When using only the FLS images for a median (as shown in Figure 2 middle), we find 5 levels of 0.085 Jy. While we accounted for the differences in overall background level, these results clearly indicate that we have some time variable aspects to the background which make it sub-optimal to construct such median background from observations about 6 months apart.
@@
To address these issues, we then processed our images through the super-background subtraction procedure developed by the SMILES (PID 1207; PI G. Rieke; Lyu et al. (2024)) team (Pérez-González et al., 2024; Alberts et al., 2024). This background subtraction starts with the stage 2 images and homogenizes the background across all images in the program taking into account the time varying behavior of the background which is particularly prominent at shorter wavelengths as here. The full details are described in Alberts et al. (2024). Figure 2 right shows the result of this super background subraction. It is clear that the residual structure in our previous simple background-subtracted image is now largely gone. In addition, this method allows for COSMOS only median since it no longer uses the fully processed images (of which there are only two), but the stage 2 images of which we have 26 dithers which is sufficient for a median background. As we show in Section 2.4 below, we also no longer find a difference in the obtained depth in FLS and COSMOS. The depth we find in FLS relative to the simple background subtraction using the FLS images alone however are comparable. This suggests that depending on the specific science applications and data available, a simple procedure as described above may be sufficient, but it should only be applied to images obtained close in time to avoid time-varying background effects.
@@ 2.3 Noise properties
To explore the noise properties of our images post background subtraction, we constructed a noise map where each pixel value equals the standard deviation in the image pixel values in a 33 pixel grid around that pixel. We performed this calculation on each image with two masks applied. First, we masked the pixels belonging to sources by applying the segmentation map. We then created a mask that removes the image edges and the residual strips, as we consider them to be sources of additional noise. This second mask is the same for each image. We perform our noise calculation on the science images after the segmentation map and noise/stripe mask have been applied. The resulting noise map for the FLS1 field is shown in Figure 3. We compared our produced noise map to the one provided with the standard pipeline products (which is derived from the per pixel coverage). The typical pixel values of these calculated noise maps are 60-70% of the MAST-provided error maps.
Figure 3 shows a wavy pattern across the whole image. We found this pattern in all our images, even when varying the number of pixels used in calculating the standard deviation, and also when running the same procedure before or after background subtraction. We also found this pattern with both the simple and super background subtracted images. We did not find this pattern when running the same procedure on one of the publicly-available CEERS images in the same filter (Finkelstein et al., 2022). The CEERS images have lower exposure times (0.82 vs. 1.85 hrs), and slightly smaller pixel scales (0.09′′ vs 0.11′′). At present, we consider this an unexplained instrumental effect that is consistent with the description of “tree rings” that have previously been observed in MIRI imaging222See https://jwst-docs.stsci.edu/jwst-calibration-pipeline-caveats/jwst-imaging-pipeline-caveats for details on the “tree ring” patterns.. This pattern is at a very low level, with the amplitude of the peaks and troughs translating to 0.5 nJy/pixel, as can also be seen in the scale color bar.
@@ 2.4 Image depth
We estimate the image depth by placing random apertures in source-free parts of the background-subtracted images. The standard deviation of the Gaussian distribution of the resultant empty aperture fluxes is the depth of the image. The above method is a simplification of the method used in Annunziatella et al. (2023), since we have fairly uniform coverage unlike that in Annunziatella et al. (2023) so noise-weighting is not critical. In Section 3.1 we describe our SExtractor photometry. For our random apertures we use a diameter equal to the median of the SExtractor Kron radii, where a Kron radius is a “reduced pseudo-radius” that defines a circular aperture containing of the flux of the source333Refer to https://sextractor.readthedocs.io/en/latest/Photom.html#automatic-aperture-flux-flux-auto and section 6 of Bertin, E. & Arnouts, S. (1996) for details.. We choose Kron apertures as they are meant to capture most of the light of our sources, which are typically spatially extended. The median Kron radius in our catalog is 0.385′′, so we adopt a diameter of 0.77′′ for our apertures. For comparison, the FWHM of the MIRI PSF at 5.6 m is 0.207′′.
We placed 500 empty apertures per field for a total of 4000 apertures. For each image, we then fitted a Gaussian to the resulting histogram of aperture flux. We found that the standard deviations of these fitted Gaussians were consistent between all the images, translating to a depth of 0.1 Jy.
As a cross-check, the median flux uncertainties for our sources based on the SExtractor photometry (Section 3.1) translate to a 5 uncertainty of 0.081 Jy. This is reasonably consistent with our image depth analysis above. Note that the random apertures method is sensitive to the adopted aperture sizes and slightly smaller apertures would bring the two values into even closer agreement. As an additional cross-check, the image depths quoted by CEERS (Yang et al., 2023), rescaled to our exposure time (since depth scales as ) translate to 0.08 Jy for our images. To be conservative, we adopt 0.1 Jy as our nominal 5 image depth.

### v2

@@ Abstract
We present an analysis of eight JWST Mid-Infrared Instrument (MIRI) 5.6 micron images with depths of . We detect 2854 sources within our combined area of 18.4 square arcminutes. We compute the MIRI number counts including an analysis of the field-to-field variation. Compared to earlier published MIRI counts, our counts have a more pronounced knee, at roughly . The location and amplitude of the counts at the knee are consistent with the Cowley et al. (2018) model predictions, although these models tend to overpredict the counts below the knee. In areas of overlap, 84 % of the MIRI sources have a counterpart in the COSMOS2020 catalog. These MIRI sources have redshifts that are mostly in the , with a tail out to . They are predominantly moderate to low stellar masses (M⊙) main sequence star-forming galaxies, suggesting that with 2 hr exposures, MIRI can reach well below at cosmic noon and reach higher mass systems out to . Nearly 70% of the COSMOS2020 sources in areas of overlap now have a data point at (rest-frame near-IR at cosmic noon) which allows for more accurate stellar population parameter estimates. Finally, we discover 31 MIRI-bright sources not present in COSMOS2020. A cross-match with IRAC channel 1 suggests that 10-20 % of these are likely lower mass (MM⊙), dusty galaxies. The rest (80–90%) are consistent with more massive but still very dusty galaxies at .
[…]
@@ 1 Introduction
Extragalactic astronomy aims to study how galaxies form and evolve across time. To do so, we build multiwavelength galaxy surveys which allow us to compare the data with spectral energy distribution (SED) models to infer redshifts and stellar population parameters (e.g. Brammer et al., 2008; Weaver et al., 2022a; Wang et al., 2024). Mid-IR data are particularly critical to this effort since they probe the rest-frame near-IR in the cosmic noon epoch and earlier which allows for much more accurate stellar masses and star formation rates (SFRs) (e.g. Elsner et al., 2008; Stefanon et al., 2015; Muzzin et al., 2009; Martis et al., 2023; La Torre et al., 2024). Indeed, over the past two decades Spitzer IRAC data – especially its first two channels, available even in the warm mission (Lacy et al., 2021; Annunziatella et al., 2023) have been a critical component of these multiwavelength surveys. With IRAC channel 1 and 2, we reached stellar masses of M at cosmic noon, i.e. well below the knee of the SMF (Elsner et al., 2008; Madau & Dickinson, 2014), but only detected the most massive galaxies () at (e.g. Stefanon et al., 2015). The IRAC channel 3, at , which critically covers the rest-frame of the stellar bump at was relatively underutilized because it was not available during the extended warm mission and was the least sensitive IRAC channel even during the cold mission.
@@
The JWST (Gardner et al., 2006), which launched on December 25, 2021, is already revolutionizing infrared astronomy with its unprecedented performance (Rigby et al., 2023). In particular, the Mid-InfraRed Instrument (MIRI; Rieke et al., 2015) allows for significantly greater sensitivity and angular resolution relative to the Spitzer Space Telescope. Recently published JWST/MIRI number counts at , and bands (Ling et al., 2022; Wu et al., 2023; Kirkpatrick et al., 2023; Stone et al., 2024) show dramatic improvement in the depth reached, even with moderate exposure times, relative to prior measurements from Spitzer and ISO. These counts have already been used in constraining galaxy and black hole evolution models (Kim et al., 2024). Number counts at the shortest MIRI wavelength () are much more scarse (Yang et al., 2023; Stone et al., 2024), but they are critical as this band samples the stellar bump at and thus is critical in testing our models of the galaxy population at cosmic noon.
@@
In addition, sampling the rest-frame near-IR at cosmic noon, the band is much less sensitive to the effects of dust obscuration even than traditionally ‘dust clean’ bands such as the band, which is rest-frame -band at the same redshifts. This insensitivity to dust obscuration is important as in the more than two decades since the discovery of the Cosmic Infrared Background (e.g. Puget et al., 1996) it has become abundantly clear that the bulk of star-formation activity at cosmic noon and beyond takes place in dust obscured galaxies (see Casey et al. 2014 for a review) which in their extreme are missed in UV/optical surveys (see e.g. Hughes et al., 1998; Sajina et al., 2006; Pope et al., 2008; Zavala et al., 2021, among many others). Recently, JWST data have further highlighted this by finding that even deep HST data miss the reddest/most dust obscured parts of the galaxy population (e.g. Labbe et al., 2023; Barrufet et al., 2023; Williams et al., 2023).
@@
In this paper, we use deep MIRI images to provide a first look at the properties of the number counts and source population based on images obtained with nearly 2 hr exposure times. This wavelength corresponds roughly to the rest-frame -band at cosmic noon (), thus probing primarily stellar mass at these critical redshifts. This paper is structured as follows. In Section 2 we present the MIRI imaging data and data reduction. In Section 3 we present the source detection and photometry and the verification of the latter. In Section 4, we show the key results of our work. These include the number counts (Section 4.2); the redshift distribution and other properties of the MIRI sources with counterparts in COSMOS2020 (Section 4.3); and lastly MIRI sources without counterparts in COSMOS2020 suggesting very red colors (Section 4.4). Throughout we adopt the AB magnitude system (Oke & Gunn, 1974). We adopt the cosmology model from the COSMOS2020 catalog (Weaver et al., 2022a), which is a standard CDM cosmology with H0 = 70 km/s/Mpc, = 0.3, and = 0.7.
@@ 2.1 MIRI imaging data
Our data come from JWST/MIRI imaging with the filter obtained as part of the GO1 program “Halfway to the Peak: A Bridge Program To Map Coeval Star Formation and Supermassive Black Hole Growth” (PIs Pope, Sajina, Yan; PID 1762; Pope et al., 2021). This program observed eight targets with the MIRI medium-resolution spectrometer (MRS) instrument: details on the targets and MRS spectra analysis are provided in Young et al. (2023). Simultaneous with the MIRI/MRS observations, we obtained MIRI imaging in fields adjacent to the MRS targets. This parallel observing means our imaging fields are effectively blank fields and therefore ideal for statistical studies such as number counts. In total, we have six fields in the First Look Survey (FLS) field and two in the COSMOS field. The FLS fields were observed in July/August 2022 and the COSMOS fields were observed in December 2022.
@@
We obtain data over the MIRI imager with an FOV of as well as the smaller MIRI coronograph with a FOV of (as seen in Figure 1). For our eight pointings, this adds up to a total area of just under 19.4 square arcminutes. The effective area we use is reduced to 18.4 square arcminutes after masking out noisy edges and bright stars. This is discussed in Section 3.2.
@@ 2.2 Data reduction
We first obtained our MIRI images from the archive already reduced using the JWST operations pipeline build 10.1 (science calibration pipeline version 1.12.5; Calibration References Data System (CRDS) version 11.17.6, context jwst_1193.pmap)111As of the time of writing, this was the latest pipeline build that affected MIRI imaging. See https://jwst-docs.stsci.edu/jwst-science-calibration-pipeline-overview/jwst-operations-pipeline-build-information. The left panel of Figure 2 shows an example of one of our images. Within these pipeline-reduced images we noted both the presence of an overall background level (the median pixel value for the FLS images was 1.23, for the COSMOS images it was 1.52). This difference in overall background level is unsurprising given that FLS is a northern field, whereas COSMOS is equatorial and as such sees a higher level of Zodiacal light, see example discussion in Sanders et al. (2007). We also noted some clear striping especially toward the bottom quarter of each image.
@@
We first attempted a simple background subtraction motivated by the fact that the striping patterns were fairly consistent among the images. We thus constructed a median background image where we used the segmentation map provided with pipeline-processed images to mask out bright sources in the fields. The segmentation map is a copy of the original image222This map is produced for each image during the source extraction step of the pipeline processing, see Section 3.1 for details regarding source extraction. where pixels are replaced either with the value 0 if they are not part of an identified source or with a positive value if they are. To avoid any remaining source signal, we added a 5 pixel “buffer” around each source (positive-value pixel) within the segmentation maps. Due to the difference in overall pixel level between the images, we rescaled each source-subtracted image to a median of one. The middle panel of Figure 2 shows an example of our image after this basic median background subtraction when we only use the FLS images for constructing the background. It is clear that this method works well in removing the vast majority of the striping though some residual structure remains due to second order differences between images. In addition, since we only have two COSMOS images we could not construct a COSMOS only median using this method. We attempted a combined median background in FLS and COSMOS; however, this led to the COSMOS fields (which have less weight in this median with 2/8 images only) to have higher noise levels. Specifically, using a single median we found 5 levels of in FLS and in COSMOS. When using only the FLS images for a median (as shown in the middle panel of Figure 2), we find 5 levels of . While we accounted for the differences in overall background level, these results clearly indicate that we have some time variable aspects to the background which make it sub-optimal to construct such median background from observations about 6 months apart.
@@
To address these issues, we then processed our images through the super-background subtraction procedure developed by the SMILES (PID 1207; PI G. Rieke; Lyu et al. (2024)) team (Pérez-González et al., 2024; Alberts et al., 2024). This background subtraction starts with the stage 2 images and homogenizes the background across all images in the program taking into account the time varying behavior of the background which is particularly prominent at shorter wavelengths as here. The full details are described in Alberts et al. (2024). The right panel of Figure 2 shows the result of this super background subtraction. It is clear that the residual structure in our previous simple background-subtracted image is now largely gone. In addition, this method allows for COSMOS only median since it no longer uses the fully processed images (of which there are only two), but the stage 2 images of which we have 26 dithers which is sufficient for a median background. As we show in Section 2.4 below, we also no longer find a difference in the obtained depth in FLS and COSMOS. The depth we find in FLS relative to the simple background subtraction using the FLS images alone however are comparable. This suggests that depending on the specific science applications and data available, a simple procedure as described above may be sufficient, but it should only be applied to images obtained close in time to avoid time-varying background effects.
@@ 2.3 Noise properties
To examine the noise characteristics of our images post background subtraction, we created a noise map. In this map, each pixel’s value represents the standard deviation in the image pixel values in a 33 pixel grid centered on that pixel. We performed this calculation on each image with two masks applied. First, we masked the pixels belonging to sources by applying the segmentation map. We then created a mask that removes the image edges and the residual strips, as we consider them to be sources of additional noise. This second mask is the same for each image. We perform our noise calculation on the science images after the segmentation map and noise/stripe mask have been applied. The resulting noise map for the FLS1 field is shown in Figure 3. We compared our produced noise map to the one provided with the standard pipeline products (which is derived from the per pixel coverage). The typical pixel values of these calculated noise maps are 60-70% of the MAST-provided error maps.
Figure 3 shows a wavy pattern across the whole image. We found this pattern in all our images, even when varying the number of pixels used in calculating the standard deviation, and also when running the same procedure before or after background subtraction. We also found this pattern with both the simple and super background subtracted images. We did not find this pattern when running the same procedure on one of the publicly-available CEERS images in the same filter (Finkelstein et al., 2022). The CEERS images have lower exposure times (0.82 vs. 1.85 hrs), and slightly smaller pixel scales (0.09′′ vs 0.11′′). At present, we consider this an unexplained instrumental effect that is consistent with the description of “tree rings” that have previously been observed in MIRI imaging333See https://jwst-docs.stsci.edu/jwst-calibration-pipeline-caveats/jwst-imaging-pipeline-caveats for details on the “tree ring” patterns.. This pattern is at a very low level, with the amplitude of the peaks and troughs translating to 0.5 nJy/pixel, as can also be seen in the scale color bar.
@@ 2.4 Image depth
We estimate the image depth by placing random apertures in source-free parts of the background-subtracted images. The standard deviation of the Gaussian distribution of the resultant empty aperture fluxes is the depth of the image. The above method is a simplification of the method used in Annunziatella et al. (2023), since we have fairly uniform coverage unlike that in Annunziatella et al. (2023) so noise-weighting is not critical. In Section 3.1 we describe our SExtractor photometry. For our random apertures we use a diameter equal to the median of the SExtractor Kron radii, where a Kron radius is a “reduced pseudo-radius” that defines a circular aperture containing of the flux of the source444Refer to https://sextractor.readthedocs.io/en/latest/Photom.html##automatic-aperture-flux-flux-auto and section 6 of Bertin, E. & Arnouts, S. (1996) for details.. We choose Kron apertures as they are meant to capture most of the light of our sources, which are typically spatially extended. The median Kron radius in our catalog is , so we adopt a diameter of for our apertures. For comparison, the FWHM of the MIRI PSF at is .
We placed 500 empty apertures per field for a total of 4000 apertures. For each image, we then fitted a Gaussian to the resulting histogram of aperture flux. We found that the standard deviations of these fitted Gaussians were consistent between all the images, translating to a depth of .
As a cross-check, the median flux uncertainties for our sources based on the SExtractor photometry (Section 3.1) translate to a 5 uncertainty of 0.081Jy. This is reasonably consistent with our image depth analysis above. Note that the random apertures method is sensitive to the adopted aperture sizes and slightly smaller apertures would bring the two values into even closer agreement. As an additional cross-check, the image depths quoted by CEERS (Yang et al., 2023), rescaled to our exposure time (since depth scales as ) translate to for our images. To be conservative, we adopt as our nominal 5 image depth.